OpenStudy (anonymous):

please help :) simplify. x/(4x) + x^2

OpenStudy (luigi0210):

I feel like I've done this problem so many times D:

OpenStudy (anonymous):

@Luigi0210 (x do you know the answer?

OpenStudy (luigi0210):

\[\frac{ x(1) }{ x(4+x) }=\frac{ 1 }{ 4+x }\]

OpenStudy (luigi0210):

I'm not giving the answer directly :P

OpenStudy (anonymous):

@Luigi0210 bleh.. confused

OpenStudy (anonymous):

@phi @ash2326 help?

OpenStudy (luigi0210):

well you have to factor out an x, and then since they have the same factors you can cancel them out

OpenStudy (luigi0210):

D: :(

OpenStudy (anonymous):

@Luigi0210 1/4??

OpenStudy (luigi0210):

@Jhannybean help me out here

OpenStudy (anonymous):

@.Sam. @Luigi0210 help me out?

sam (.sam.):

\[\frac{x}{4x+x^2}\] Factor x from denominator \[=\frac{x}{x(4+x)}\] cancel x \[=\frac{\cancel x}{\cancel x(4+x)} \\ \\ =\frac{1}{4+x}\]

OpenStudy (luigi0210):

I tried explaining it :/

OpenStudy (anonymous):

@.Sam. i love you. can you help me out with this? simplify. (-14x^3)(/x^3-5x^4)

OpenStudy (anonymous):

@Luigi0210 aw, you helped too. :)

OpenStudy (luigi0210):

:(

sam (.sam.):

Same thing, factor x^3 from denominator then cancel it

OpenStudy (anonymous):

@.Sam. -14x/1x-5?

OpenStudy (phi):

that looks like one of the choices, i.e. a pure guess.

sam (.sam.):

should be -14/1-5x

OpenStudy (anonymous):

@.Sam. ohh

sam (.sam.):

\[\frac{-14x^3}{x^3-5x^4}=\frac{-14\cancel{x^3}}{\cancel{x^3}(1-5x)}=\frac{-14}{1-5x}\]

OpenStudy (anonymous):

@.Sam. ohhh, I see!

OpenStudy (anonymous):

@.Sam. so for x+7/x^2+4x-2, would it be 1/x-7??

sam (.sam.):

look at the denominator, looks like it can't be factored

OpenStudy (anonymous):

@.Sam. you're right.. ach, i feel so stupid! :/

OpenStudy (anonymous):

@.Sam. wait, is it 1/x-3?

sam (.sam.):

no, if you cant factor, the answer will just be x+7/x^2+4x-2

OpenStudy (anonymous):

@.Sam. but, that's not an option

sam (.sam.):

maybe you wrote the question wrongly

OpenStudy (anonymous):

@.Sam. no, that's the question

sam (.sam.):

\[\Huge \frac{x+7}{x^2+4x-2}\]